OpenStudy (anonymous):

*** I will give you a medal to answer *** Find the least common multiple of x³ – x² + x – 1 and x² – 1 . Write the answer in factored form.

OpenStudy (anonymous):

\[x^3-x^2+x-1=(x^2+1)(x-1)\] and \(x^2-1=(x+1)(x-1)\)

OpenStudy (anonymous):

so the lcm is \[(x-1)(x+1)(x^2+1)\]

OpenStudy (anonymous):

thank you (: